OpenStudy (anonymous):

simplify -8x+2x-x+15x

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

and fan maybe?

OpenStudy (anonymous):

-8x + 2x - x + 15x -8x -x = -9x 15 x + 2x = 17x 17x - 9x = 8x the final answer is 8x